# Erasing Conceptual Knowledge from Language Models

### v1

@@ 5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) Zou et al. (2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the following example:
We provide more details about GCG attack and baseline robustness details in Appendix C.
@@ 5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by Li et al. (2024).
@@ 5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the task of erasing knowledge related to the Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
| Method | Innocence () | Specificity () | Seamlessness () |
@@
| Original Model (Touvron et al., 2023) | 66.4 | 47.0 | 3.6 |
| RMU (Li et al., 2024) | 51.0 | 44.2 | 3.7 |
| WHP (Eldan & Russinovich, 2023) | 58.6 | 43.1 | 3.4 |
| ELM (Ours) | 38.3 | 45.3 | 3.4 |
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to Lynch et al. (2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge. RMU (Li et al., 2024) proved to be ineffective in erasing with our initial large hyperparameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
@@
While ELM demonstrates strong performance in targeted concept removal, a few limitations warrant further study. We observe a noticable degradation in MMLU scores for nearby safer concepts. This challenge calls for more fine-grained erasure techniques and optimization constraints to minimize unintended knowledge loss. Second, while ELM generates fluent text when prompted with erased concepts, in a few cases, outputs lack meaningful content or contextual relevance. This highlights the difficulty of maintaining semantic coherence while navigating around removed knowledge, suggesting the need for more sophisticated language modeling approaches during erasure. The effectiveness of ELM in erasing complex, interconnected concepts remains an open question. Language models encode information in intricate ways, with subtle dependencies between semantic domains. Completely isolating and removing specific concepts without affecting related knowledge presents an ongoing challenge, necessitating more nuanced erasure techniques that can handle concept interdependencies. Finally, while our evaluation framework comprehensively assesses erasure quality, it may not capture all aspects of model behavior affected by the process. Future work should focus on continuing the development of more rigorous evaluation metrics and exploring the wider effects of concept erasure on model performance and generalization capabilities.
@@ 7 Conclusion
This work introduces a framework for evaluating concept erasure in language models, centered on three critical criteria: innocence, seamlessness, and specificity. Our proposed Erasure of Language Memory (ELM) method emerges as a direct consequence of this framework. ELM uses low-rank model updates to selectively alter output distributions for erased concepts while preserving overall model capabilities and fluency. The erasing objective effectively eliminates targeted concept knowledge, as evidenced by near-random performance on multiple-choice questions related to erased topics. Simultaneously, the retention and fluency objectives ensure coherent text generation and preserved performance on unrelated tasks. Furthermore, ELM demonstrates robust resistance to adversarial attacks which underscores the effectiveness of our approach in creating models that are not only “innocent” of erased knowledge but also maintain their functionality and coherence. In addition to providing a practical solution for concept erasure, we have established a foundation for more comprehensive evaluation of knowledge erasure in language models.
@@ Acknowledgements
Our methods are available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
@@ Appendix A Details on metrics
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP Li et al. (2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
@@ B.1 WMDP Results
We directly download the best Zephyr-7B RMU model from the WMDP authors (https://huggingface.co/cais/Zephyr_RMU) for testing.
@@ Appendix D Hyperparameter Analysis
Appendix E Conditional Fluency Training
@@
Appendix F Progression of ELM Training

### v2

@@ 5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) (Zou et al., 2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below. We provide GCG attack details in Appendix C. We also show ELM’s robustness to more attacks in Appendix H.
@@ 5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by (Li et al., 2024).
@@ 5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to (Lynch et al., 2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge. RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
| Method | Innocence () | Specificity () | Seamless () |
@@
| Original | 66.4 | 47.0 | 3.6 |
| RMU | 51.0 | 44.2 | 3.7 |
| WHP | 58.6 | 43.1 | 3.4 |
| ELM | 38.3 | 45.3 | 3.4 |
@@
While ELM effectively removes targeted concepts through introspective classification, several limitations merit investigation. The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge. Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence - suggesting that our probability modification may be overly aggressive in some cases. The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base. Further work is needed to develop more granular techniques for selective knowledge modification while preserving complex conceptual inter-dependencies.
@@ 7 Conclusion
This work reframes the challenge of machine unlearning for large language models, shifting from traditional sample-based approaches to concept-oriented unlearning through introspective classification. Our proposed Erasure of Language Memory (ELM) method demonstrates that effective concept unlearning requires modifying the model’s output distribution based on its own ability to recognize and evaluate knowledge. By using low-rank model updates guided by the model’s introspective classification, ELM achieves targeted concept removal while preserving the model’s broader capabilities. Our experiments show that this approach overcomes limitations of previous methods like gradient ascent or representation disruption, as evidenced by near-random performance on multiple-choice questions related to erased concepts while maintaining accuracy on other tasks. Furthermore, ELM’s resistance to adversarial attacks validates our hypothesis that concept unlearning should leverage the model’s own understanding of its knowledge. In addition to providing a practical solution for concept erasure, we have established a foundation for more comprehensive evaluation of knowledge erasure in language models.
@@ Acknowledgements
ELM is available as open-source code. Source code, trained models, and data sets for reproducing our results can be found at elm.baulab.info and at https://github.com/rohitgandikota/erasing-llm.
@@ Appendix A Details on metrics
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP (Li et al., 2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
@@ B.1 WMDP Results
We directly download the best Zephyr-7B RMU model from the WMDP authors (https://huggingface.co/cais/Zephyr_RMU) for testing. For Mistral, we run a hyperparameter sweep over , layer indices 3,4,5, 4,5,6, and 5,6,7, and learning rates . We select runs with the lowest possible WMDP accuracies that don’t completely destroy MMLU accuracy. For Mistral, this is and lr= at layers 5,6,7. We sweep across the same hyperparameters for Llama-3-8B. Llama-3-8B-Instruct uses the best hyperparameters found in the base model sweep. The runs shown in Table 1 have and lr= at layers 4,5,6. All runs had a steering coefficient of 6.5.
@@
For Mistral, we run a hyperparameter sweep over , , and learning rates . We selected the run that has the lowest possible WMDP accuracies without destroying MMLU accuracy. This run, shown in Table 1, has the parameters , , lr=.
We run a sweep over the same hyperparameters for Llama-3-8B, and use the best runs from the base model to decide hyperparameters for Llama-3-8B-Instruct. The runs shown in Table 1 had , , lr=.
@@ Appendix D Hyperparameter Analysis
Appendix E Low-Rank vs Full Finetuning
We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints. In Table 4, we show the performance of ELM on Zephyr-7B when editing with full finetuning and low-rank model editing. Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model. Where as low-rank model editing preserves the specificity while being effective at erasure.
| Method | Innocence () | Specificity () | ||
| Bio | Cyber | MMLU | MT-Bench | |
| Original | 64.4 | 44.3 | 58.5 | 7.3 |
| ELM - Full | 25.4 | 27.1 | 45.2 | 3.4 |
| ELM - LoRA | 29.7 | 27.2 | 56.6 | 7.1 |
Appendix F Conditional Fluency Training
@@
Appendix G Progression of ELM Training
@@
Appendix H Adversarial Attacks
We also attack ELM with BEAST (Sadasivan et al., 2024), a fast adversarial prompt based attack on LLMs. BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response. We find that BEAST is unable to extract erased information from ELM:
Additionally - finetuning attack where we train ELM model autoregressively on the original forget dataset. We find that the resulting attacked model brings back the knowledge slightly (Bio: 29.7% to 42.2%; Cyber: 27.2% to 29.4%) but not to the original level of 64.4% Bio and 44.3% Cyber. ELM models can be retrained to bring back erased knowledge, but it is harder.